# From Intention To Implementation: Automating Biomedical Research via LLMs

### v2

@@ 3.2 Search
(1) Query Generation: Searching directly with the user input can yield imprecise results for two reasons. Firstly, the user describes the research objective and conditions in natural language, while most databases use Boolean query logic, making it difficult to retrieve results that completely match the lengthy user input. Secondly, the user intention is provided with key high-level concepts, while the literature may use a more detailed description or synonyms, making it challenging to obtain all relevant materials. Thus, query rewriting is crucial (Clark, 2013). However, manually crafting effective queries costs a lot of expertise, labor, and time (Scells et al., 2020). To address this, the Search module in BioResearcher employs an LLM-based query generator agent to create Boolean queries based on the user input. As displayed in Figure 3, the query generator extracts keywords to improve the retrieval accuracy. It also performs synonym expansion (e.g., expanding ”Single-cell sequencing” with ”scRNA-seq”) to improve the retrieval recall. These structured queries allow the module to interpret the research objective effectively and focus on the most relevant materials.
Figure 3: An example of query generation. The query generator generates five Boolean queries for each user input.
@@
(3) Filtration: To retain only the most relevant and useful materials for further stages, our system employs an LLM-based filter agent to filter the returned research articles and datasets. For the research articles, we define a set of criteria detailed in Table 1. The filter examines the titles and abstracts of each article to determine their potential contribution to the research objectives. Each article receives a helpfulness score ranging from one to five. Articles scoring above four are downloaded and forwarded to the Literature Processing module. For the datasets, the filter assesses its usefulness based on the meta-data (i.e., the attached online descriptions) of datasets and assigns a binary usability score. The useful datasets’ descriptions are forwarded to the Experimental Design module.
Table 1: Scoring criteria for the helpfulness of papers.
Rating Description
1 Not Helpful at All:
2 Slightly Helpful:
3 Moderately Helpful:
4 Very Helpful:
5 Extremely Helpful:
@@ 3.3 Literature Processing
Literature comprehension can be challenging for both researchers and LLMs due to the massive, lengthy, unstructured, and logically complex nature of research papers. To address this, we introduce the Literature Processing module, which streamlines comprehension and aids in experimental design. This module first standardizes research papers into structured experimental reports and systematically analyzes them. It then interacts with the Search module to assess the usability of datasets mentioned in these reports. The module operates in two main phases: report generation and report analysis. Figure 4 demonstrates the module in action.
Figure 4: An example of Literature Processing. (a) illustrates the workflow for Report Generation, and (b) presents an example of Report Analysis. An LLM-based reviewer agent is involved in both processes.
@@
To mitigate potential performance degradation caused by excessively long input contextsBai et al. (2024), the report is divided into sections corresponding to the first-level headings, allowing parallel processing and enhancing efficiency.
@@
Drawing inspiration from the Chain of Thought (COT) framework (Wei et al., 2022), which emphasizes step-by-step reasoning, we recognize that analyzing reports in relation to research objectives is a critical step in designing new experiments. Consequently, we introduce a report analysis process following report generation. Specifically, an LLM-based analyst agent evaluates each section of the experimental report for its referability, considering the research objective, conditions, and requirements. The analyst also provides suggestions for references and modifications, akin to proposing innovations on existing methods. Additionally, this module interfaces with the Search module to identify usable datasets within the report, using regular expressions to extract dataset IDs for retrieval from the database while bypassing the query generation step. As mentioned above, the usefulness of each dataset is determined based on its description, and useful datasets are then integrated with previously collected ones.
@@ 3.4 Experimental Design
Scientific findings in biomedical research frequently face reproducibility issues, wasting resources and time while undermining the credibility of scientific outcomes (Begley & Ioannidis, 2015). A well-designed experimental protocol is crucial for obtaining reliable results and optimizing resource use (Larijani et al., 2020). Therefore, we develop the Experimental Design module, which uses an LLM-based experiment designer agent to create scientific and reproducible protocols. We propose a hierarchical learning approach to ensure the rationality of the logical structure of the generated experimental protocols. This method employs the Retrieval-Augmented Generation (RAG) technique aided by the analysis (i.e., using relevant reports’ sections with high referability as reference materials), enabling the model to subsequently learn first-level headings independently, then outlines, and then experimental details from the reorganized reports. The design process is structured into three essential steps, as demonstrated in Figure 5.
Figure 5: Demonstration of experimental design. The design process consists of three steps: (1) First-Level Headings Design, (2) Outline Generation, and (3) Generation of Implementation Details. An LLM-based reviewer agent is involved in each step.
@@ 4.1 Protocol Evaluation
Completeness
Completeness assesses how thoroughly each section of the protocol is described, considering the necessary steps that should be added to achieve the design purpose of this part. The formula for completeness is given by:
@@
Level of Detail
The level of detail measures the degree to which a protocol provides sufficient information for each step, ranging from 0 (no detail) to 1 (fully detailed).
Correctness
Correctness assesses the proportion of protocol steps that are free from factual errors. Our analysis reveals that protocols with shorter steps tend to exhibit higher correctness scores, as the probability of factual errors increases with longer and more detailed steps. Drawing inspiration from the BLEU (Papineni et al., 2002) metric in machine translation, we also introduce a brevity penalty (BP) for shorter steps. The BP is constrained to a minimum of 0.5, and the formula is as follows:
$\text{BP}=\begin{cases}1,&\text{if }l_{steps}>L\\ max(e^{(1-\frac{L}{l_{steps}})},~{}0.5),&\text{if }l_{steps}\leq L\end{cases}$ (2)
@@
Logical Soundness
Logical soundness evaluates the proportion of steps that are placed in a reasonable order within a protocol. Reasonable steps are those that are logically ordered and appropriately positioned. The formula for Logical Soundness is given by:
@@
Structural Soundness
Structural soundness evaluates the logical coherence and organizational integrity of a protocol’s overall framework, with scores from 0 (completely unsound) to 1 (perfectly sound).
@@ 5.1 Experiment Setup
We evaluate BioResearcher by comparing it with three well-known agent systems. (1) ReAct (Yao et al., 2023), which integrates reasoning and action within LLMs to effectively manage complex reasoning and decision-making tasks. (2) Plan-and-Execute (Wang et al., 2023c), which utilizes an iterative framework to accomplish tasks through sequential planning and execution. (3) RAG, which employs a naive Retrieval-Augmented Generation (RAG) module to search for relevant content and generates the answer in a single step. Section A provides a detailed description of the implementation of these baseline systems.
@@ 5.2 Performance of End-To-End Automation
We collect eight ongoing research objectives from a biomedical laboratory to ensure that no published work has addressed these research objectives, as detailed in Table 9. Each objective is respectively processed for three runs to minimize randomness. We evaluate three baseline systems and our system for designing and executing experiments for these objectives. We equip the React and Plan-and-Execute systems with four tools: (1) a search tool utilizing the NCBI API  [^4] https://www.ncbi.nlm.nih.gov/home/develop/api/ to retrieve descriptions of relevant papers, (2) a download tool for acquiring these papers and storing them in a chunked and vectorized format, (3) a search tool using the NCBI API to obtain descriptions of relevant datasets and storing them in a chunked and vectorized format, and (4) a search tool that extracts pertinent content from the resulting vector database.
@@
RAG 0.392 0.765 0.470 0.991 0.952 3.570 1.286 9.167
ReAct 0.361 0.617 0.487 0.972 0.950 3.387 1.237 5.792
Plan and Execute 0.366 0.617 0.487 1.000 0.935 3.405 1.194 6.375
BioResearcher $\textbf{0.612}_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow\textbf{0.220}}}$ $\textbf{0.902}_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow\textbf{0.137}}}$ $\textbf{0.944}_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow\textbf{0.457}}}$ $0.987_{\downarrow 0.013}$ $0.910_{\downarrow 0.042}$ $\textbf{4.355}_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow\textbf{0.785}}}$ 7.327 33.958
5.2.1 Results
Table 3 presents the average scores for protocols generated by different systems, including Completeness, Level of Detail, Correctness, Logical Soundness, and Structural Soundness. We also calculate the overall performance, the average number of sentences per step ($l_{\text{steps}}$), and the average number of steps per protocol ($n_{\text{total steps}}$). (1) The results highlight BioResearcher’s exceptional performance in Completeness, Level of Detail, and Correctness, surpassing the best baseline by 0.220 (56.1%), 0.137 (17.9%), and 0.457 (93.8%), respectively. (2) Our system is comparable to the best baseline in Logical Soundness and Structural Soundness, exceeding 0.9, consistently maintaining a high performance in these aspects. (3) Furthermore, the protocols generated by BioResearcher have significantly more sentences per step and steps per protocol, 5.9×5.9$\times$5.9 × and 4.8×4.8$\times$4.8 × greater than the average performance of different baselines, respectively, indicating the generation of more detailed and comprehensive protocols. (4) The comparative analysis of the three baselines indicates that RAG outperforms the other two iterative agent systems. This superiority can be attributed to the fact that, in the long-context environment characterized by iteration and lengthy retrieved information, the summarization and integration capabilities of the latter two systems degrade, resulting in the loss of substantial amounts of valuable information in the final generated protocol. In contrast, we design tailored workflows to effectively integrate results from multiple modules and steps, thereby preserving critical information throughout the integration process. A case comparison of protocols from the four systems is illustrated in Figure 10.
@@ 5.2.1 Results
5.2.2 The Quality of LLM Evaluation
To testify whether the evaluation gained by the LLM judge is truthful, we engage both human experts and the LLM to evaluate 18 protocols using consistent criteria focused on a single research objective. By concentrating on a single objective, we reduce the evaluation difficulty for human experts, thereby enhancing the accuracy of their assessments due to their deeper understanding of the specific context. Subsequently, human experts review the LLM-generated evaluations to verify their precision. The premise of this experiment is based on the assumption that while human experts might not identify all issues, any issues they do identify are accurate, and they can reliably judge the validity of issues identified by the model.
Table 4: Performance of LLM and human experts’ evaluation.
Metrics (%) Completeness Detail Correctness Logical Soundness Structure Avg.
LLM’s Precision 81.4 100.0 74.0 83.3 79.6 83.6
LLM’s Recall 100.0 - 51.5 74.4 - 75.3
Experts’ Recall 44.2 - 70.7 75.6 - 63.5
We use two commonly adopted metrics for classification: Precision and Recall. In this context, precision is defined as the ratio of true positive results to the total number of positive predictions made by the LLM, where a true positive is an issue identified by the LLM and confirmed as valid by human experts. Recall is defined as the ratio of true positive results to the total number of actual positive cases. The total actual positive cases are the union of two sets: (1) issues identified by the LLM and validated by human experts, and (2) issues independently identified by human experts.
The results are summarized in Table 4. Specifically, the LLM reviewer achieves an average precision of 83.6% across five dimensions, with a remarkable precision of 100.0% in the level of detail, indicating the reliability of LLM assessments. Additionally, we find that the LLM’s average recall in completeness, correctness, and logical soundness exceeds that of human experts by 11.8%. Notably, in terms of completeness, the LLM successfully identifies all sections of protocols with missing steps. This suggests that the LLM has a superior ability to identify errors and omissions when evaluating protocols step-by-step compared to human experts. In summary, the above results verify that LLMs can evaluate the automation performance more thoroughly, accurately, and quickly.
@@ 5.3.1 The Effect of Generated Queries
We conduct a comparative experiment to assess the effectiveness of LLM-generated queries. An LLM and three human participants independently generate five queries for each user input, which specifies a research objective, conditions, and requirements. The evaluation metric is the number of relevant papers or datasets retained after retrieval and filtering. The LLM generates queries three times, with the results averaged, and a similar average is calculated across the three human participants. This experiment, covering ten research objectives listed in Table 10, presents its results in Figure 8.

### v4

@@ 3.2 Search
(1) Query Generation: Searching directly with the user input can yield imprecise results for two reasons. Firstly, the user describes the research objective and conditions in natural language, while most databases use Boolean query logic, making it difficult to retrieve results that completely match the lengthy user input. Secondly, the user intention is provided with key high-level concepts, while the literature may use a more detailed description or synonyms, making it challenging to obtain all relevant materials. Thus, query rewriting is crucial [63]. However, manually crafting effective queries costs a lot of expertise, labor, and time [64]. To address this, the Search module in BioResearcher employs an LLM-based query generator agent to create Boolean queries based on the user input. As displayed in Figure 3, the query generator extracts keywords to improve the retrieval accuracy. It also performs synonym expansion (e.g., expanding ”Single-cell sequencing” with ”scRNA-seq”) to improve the retrieval recall. These structured queries allow the module to interpret the research objective effectively and focus on the most relevant materials.
Figure 3: An example of query generation.
@@
(3) Filtration: To retain only the most relevant and useful materials for further stages, our system employs an LLM-based filter agent to filter the returned research articles and datasets. For the research articles, we define a set of criteria detailed in Appendix D (Table D1). The filter examines the titles and abstracts of each article to determine their potential contribution to the research objectives. Each article receives a helpfulness score ranging from one to five. Articles scoring above four are downloaded and forwarded to the Literature Processing module. For the datasets, the filter assesses its usefulness based on the metadata (i.e., the attached online descriptions) of datasets and assigns a binary usability score. The useful datasets’ descriptions are forwarded to the Experimental Design module.
@@ 3.3 Literature Processing
Literature comprehension can be challenging for researchers and LLMs because the research papers are massive, lengthy, and unstructured with complex logic. To streamline literature comprehension, enhance comprehension, and provide valuable references for experimental design, we introduce the Literature Processing module. This module first standardizes research papers into highly structured experimental reports and analyzes them systematically. It then interacts with the Search module to identify the usability of the datasets mentioned in the reports. Thus, this module operates through two primary phases: report generation and report analysis. Figure 4 illustrates an example of the module in operation.
Figure 4: An example of Literature Processing. (a) illustrates the workflow for Report Generation, while (b) presents an example of Report Analysis.
@@
To mitigate potential performance degradation caused by excessively long input contexts [65], the report is divided into sections corresponding to the first-level headings, allowing parallel processing and enhancing efficiency.
@@
Drawing inspiration from the Chain of Thought (COT) framework [66], which emphasizes step-by-step reasoning, we recognize that analyzing reports in relation to research objectives is a critical step in designing new experiments. Consequently, we introduce a report analysis process following report generation. Specifically, an LLM-based analyst agent evaluates each section of the experimental report for its referability, considering the research objective, conditions, and requirements. The analyst also provides suggestions for references and modifications, akin to proposing innovations on existing methods. Additionally, this module interfaces with the Search module to identify usable datasets within the report, using regular expressions to extract dataset IDs for retrieval from the database while bypassing the query generation step. As mentioned above, the usefulness of each dataset is determined based on its description, and useful datasets are then integrated with previously collected ones.
@@ 3.4 Experimental Design
Scientific findings in biomedical research frequently face reproducibility issues, wasting resources and time while undermining the credibility of scientific outcomes [67]. A well-designed experimental protocol is crucial for obtaining reliable results and optimizing resource use [68]. Therefore, we develop the Experimental Design module, which uses an LLM-based experiment designer agent to create scientific and reproducible protocols. We propose a hierarchical learning approach to ensure the rationality of the logical structure of the generated experimental protocols. This method employs the Retrieval-Augmented Generation (RAG) technique aided by the analysis (i.e., using relevant reports’ sections with high referability as reference materials), enabling the model to subsequently learn first-level headings independently, then outlines, and then experimental details from the reorganized reports. The design process is structured into three essential steps, as demonstrated in Figure 5.
Figure 5: Demonstration of experimental design.
@@ 4.1 Protocol Evaluation
  • Completeness: Completeness assesses how thoroughly each section of the protocol is described, considering the necessary steps that should be added to achieve the design purpose of this part. The formula for completeness is given by:
@@
  • Level of Detail: The level of detail measures the degree to which a protocol provides sufficient information for each step, ranging from 0 (no detail) to 1 (fully detailed).
  • Correctness: Correctness assesses the proportion of protocol steps that are free from factual errors. Our analysis reveals that protocols with shorter steps tend to exhibit higher correctness scores, as the probability of factual errors increases with longer and more detailed steps. Drawing inspiration from the BLEU [25] metric in machine translation, we also introduce a brevity penalty (BP) for shorter steps. The BP is constrained to a minimum of 0.5, and the formula is as follows:
    $\text{BP}=\begin{cases}1,&\text{if }l_{steps}>L,\\ max(e^{(1-\frac{L}{l_{steps}})},~{}0.5),&\text{if }l_{steps}\leq L,\end{cases}$ (2)
@@
  • Logical Soundness: Logical soundness evaluates the proportion of steps that are placed in a reasonable order within a protocol. Reasonable steps are those that are logically ordered and appropriately positioned. The formula for Logical Soundness is given by:
@@
  • Structural Soundness: Structural soundness evaluates the logical coherence and organizational integrity of a protocol’s overall framework, with scores from 0 (completely unsound) to 1 (perfectly sound).
@@ 5.1 Experiment Setup
We evaluate BioResearcher by comparing it with three well-known agent systems. (1) ReAct [69], which integrates reasoning and action within LLMs to effectively manage complex reasoning and decision-making tasks. (2) Plan-and-Execute [27], which utilizes an iterative framework to accomplish tasks through sequential planning and execution. (3) RAG, which employs a naive Retrieval-Augmented Generation (RAG) module to search for relevant content and generates the answer in a single step. Appendix B provides a detailed description of the implementation of these baseline systems.
@@ 5.2 Performance of End-To-End Automation
We collect eight ongoing research objectives from a biomedical laboratory to ensure that no published work has addressed these research objectives, as detailed in Appendix C (Table C1). Each objective is processed for three runs to minimize randomness. We evaluate three baseline systems and our system for designing and executing experiments for these objectives. We equip the React and Plan-and-Execute systems with four tools: (1) a search tool utilizing the NCBI API  [^5] https://www.ncbi.nlm.nih.gov/home/develop/api/ to retrieve descriptions of relevant papers, (2) a download tool for acquiring these papers and storing them in a chunked and vectorized format, (3) a search tool using the NCBI API to obtain descriptions of relevant datasets and storing them in a chunked and vectorized format, and (4) a search tool that extracts pertinent content from the resulting vector database.
@@
RAG 0.405 0.687 0.483 0.973 0.908 3.456 1.286 9.167
ReAct 0.364 0.577 0.484 0.963 0.897 3.285 1.237 5.792
Plan and Execute 0.380 0.587 0.483 0.965 0.900 3.314 1.194 6.375
BioResearcher $\textbf{0.659}_{\uparrow\textbf{0.254}}$ $\textbf{0.893}_{\uparrow\textbf{0.206}}$ $\textbf{0.895}_{{\uparrow\textbf{0.411}}}$ $0.953_{\downarrow 0.020}$ $0.891_{\downarrow 0.017}$ $\textbf{4.292}_{{\uparrow\textbf{0.836}}}$ 7.327 33.958
5.2.1 Performance by Automatic Evaluation
Table 2 presents the average scores for protocols generated by different systems, including Completeness, Level of Detail, Correctness, Logical Soundness, and Structural Soundness. To minimize single-model evaluation bias, we employed four LLMs (GPT-4o, O3-mini-2025-01-31, Gemini-2.0-Flash, and DeepSeek-V3) for assessment, reporting their average scores across five distinct metrics. The detailed scoring of each model can be found in Appendix F (Table F1). We also calculate the overall performance, the average number of sentences per step ($l_{\text{steps}}$), and the average number of steps per protocol ($n_{\text{total steps}}$). (1) The results highlight BioResearcher’s exceptional performance in Completeness, Level of Detail, and Correctness, surpassing the best baseline by 0.254 (62.7%), 0.206 (30.0%), and 0.411 (84.9%), respectively. (2) Our system is comparable to the best baseline in Logical Soundness and Structural Soundness, exceeding 0.89, consistently maintaining a high performance in these aspects. (3) Furthermore, the protocols generated by BioResearcher have significantly more sentences per step and steps per protocol, 5.9×5.9$\times$5.9 × and 4.8×4.8$\times$4.8 × greater than the average performance of different baselines, respectively, indicating the generation of more detailed and comprehensive protocols. (4) The comparative analysis of the three baselines indicates that RAG outperforms the other two iterative agent systems. This superiority can be attributed to the fact that, in the long-context environment characterized by iteration and lengthy retrieved information, the summarization and integration capabilities of the latter two systems degrade, resulting in the loss of substantial amounts of valuable information in the final generated protocol. In contrast, we design tailored workflows to effectively integrate results from multiple modules and steps, thereby preserving critical information throughout the integration process. A case comparison of protocols from the four systems is illustrated in Appendix F (Figure F1).
@@ 5.2.1 Performance by Automatic Evaluation
5.2.2 Quality of Automatic Evaluation
To validate the reliability of evaluations conducted by the LLM judge, we engage three domain experts to independently assess the LLM judge’s evaluation outcomes. Given the time-consuming and labor-intensive nature of manual evaluation, we employed only GPT-4o, the best LLM currently, as the foundation LLM in BioResearcher to conduct multiple sampling generations on one research objective, resulting in the generation of 18 protocols used in this experiment. This focused approach also reduces evaluation complexity while enhancing assessment accuracy through deeper contextual understanding by experts.
Table 3: Performance of the LLM judge, as evaluated by human experts. Human consistency is measured using Fleiss’ kappa.
Completeness Detail Structure Logical Soundness Correctness Avg.
LLM’s accuracy (%) 91.1 88.9 75.9 75.0 71.3 80.4
Human consistency 0.66 0.92 0.90 0.88 0.95 0.86
We engage three domain experts to systematically review the LLM judge’s evaluation rationales and results for each protocol across five dimensions. This process involves verifying the factual accuracy and logical credibility of the LLM’s outputs. Expert judgments are aggregated through majority voting to determine the validity of each step-level evaluation. The final analysis quantifies the LLM’s assessment accuracy across five dimensions.
Results are shown in Table 3, demonstrating that the LLM achieves an overall accuracy of 80.4% in protocol evaluations, with exceptional performance in Completeness (91.1%). These metrics confirm the model’s capacity to deliver comprehensive and efficient automation assessments. We also calculate the inter-expert consistency using Fleiss’ kappa to ensure evaluation reliability. substantial agreement is observed across all dimensions: Detail ($\kappa=0.92$), Structure ($\kappa=0.90$) and Correctness ($\kappa=0.95$) reaches near-perfect consensus, while other dimensions maintain $\kappa>0.65$. The p-values for consistency across the three dimensions are uniformly below 0.05. This agreement underscores the methodological rigor and reproducibility of the evaluation framework.
5.2.3 Performance by Manual Evaluation
To further validate the quality of the protocols generated by BioResearcher, we engage three domain experts to independently assess the 18 original protocols generated on one research objective, same as in Section 5.2.2, using consistent criteria in Section 4.1. Concentrating on one research objective mitigates assessment complexity and enhances accuracy through deeper contextual understanding by the experts.
Table 4: Quality of experimental protocols evaluated by human experts. Consistency for the Correctness metric is measured using Fleiss’ kappa, while consistency for the Detail and Structure metrics is assessed using Kendall’s W.
Correctness Detail Structure
Human Experts 0.85 0.87 0.87
Human Consistency 0.85 0.68 0.61
We employ experts to evaluate these protocols across three dimensions: Correctness, Detail, and Structure, following the criteria in Section 4.1. Specifically, the experts systematically evaluate each step of the protocols, determining whether it is correct or not to derive the $n_{cs}$ in equation 3 for calculating the correctness. For the remaining two metrics, the experts assign a direct score to each protocol. We exclude Completeness and Logical Soundness due to their susceptibility to subjective interpretation, as different researchers may design divergent yet valid experimental approaches. We then analyze inter-rater reliability using appropriate statistical measures. For Correctness, which quantifies the proportion of correct experimental steps in a protocol, we employ Fleiss’ kappa. For Detail and Structure (ordinal scoring), we employ Kendall’s W. As shown in Table 4, the results in three dimensions are all over 85%, and all consistency metrics exceed 0.6, with Correctness over 0.8. The p-values for consistency across the three dimensions are uniformly below 0.05. These results confirm strong consensus among evaluators across all dimensions.
@@ 5.3.1 The Effect of Generated Queries
We conduct a comparative experiment to assess the effectiveness of LLM-generated queries. An LLM and three human participants independently generate five queries for each user input, which specifies a research objective, conditions, and requirements. The evaluation metric is the number of relevant papers or datasets retained after retrieval and filtering. The LLM generates queries three times, with the results averaged, and a similar average is calculated across the three human participants. This experiment, covering ten research objectives listed in Appendix C (Table C2), presents its results in Figure 8.